OpenStudy (anonymous):

find the first 9th term of power series of sin x

OpenStudy (amistre64):

derviables :)

OpenStudy (anonymous):

\[ \sin(x) = \sum^\infty_{k=0} (-1)^k\frac{x^{2k+1}}{(2k+1)!} \] You should be able to write out the first 9 terms.

OpenStudy (anonymous):

x-x^3/6+x^5/120-x^7/5040+x^9/362880+O(x^10)

OpenStudy (anonymous):

But thats only 5 terms..

OpenStudy (anonymous):

\[ \sin(x)=x-{\frac {1}{6}}{x}^{3}+{\frac {1}{120}}{x}^{5}-{\frac {1}{5040}}{x}^ {7}+{\frac {1}{362880}}{x}^{9}-{\frac {1}{39916800}}{x}^{11}+{\frac {1 }{6227020800}}{x}^{13}-{\frac {1}{1307674368000}}{x}^{15}+{\frac {1}{ 355687428096000}}{x}^{17}+O \left( {x}^{19} \right) )\]